### v1

@@ Results
To carry out our analyses, we compute the mean squared displacement (MSD) of the particle positions as a function of time (see Methods for details) for simulations with different densities and setups, to compare their corresponding dynamics based on the resulting curves. We change the density by varying the disk diameter $l_{0}$ while keeping constant the number of agents and the size of the arena. We study two types of boundary conditions, analyzing simulations in arenas with smooth or rough circular confinements, while focusing on anisotropic damping. Finally, we consider two scenarios for the off-centered rotation distance: a small $R=0.03$ case and a large $R=0.3$ case, corresponding to weak and strong explicit alignment, respectively. Note that $R$ is defined here as independent of disk diameter, so changing the value of $l_{0}$ will also change the relative off-centered rotation distance with respect to its radius. We thus find that, for $R=0.03$ and $l_{0}=0.8,1.0,1.2$, we have $R/(l_{0}/2)=0.075,0.06,0.05<<1$, all representing weak alignment. For $R=0.3$, we have $R_{r}=0.75,0.6,0.5$, which correspond instead to strong alignment.
@@
Figure 7 displays snapshots and the MSD for simulations in a smooth boundary arena, with different densities given by low ($\phi=0.580$) and high ($\phi=0.907$, $\phi=1.306$) packing fractions. The corresponding temporal dynamics are displayed as videos in the Supplemental Material. Supplementary Movies 9-11 correspond to the snapshots in Fig. 7a-c and Supplementary Movies 12-14 corresponds to the snapshots in Fig. 7e-g. For the small $R=0.03$ case, Panel (a) shows that, at low densities, we find clear milling motion around an empty center in a state characterized by order parameters $P=0.012$ and $M=0.997$. At higher densities, this milling state is suppressed as the central region is filled with disks and the order parameters become $P=0.387$ and $M=0.753$ in Panel (b) and $P=0.308$, $M=0.821$ in Panel (c), which corresponds more to an aligned state displaying localized rotation. On the other hand, for the large $R=0.3$ case in Panels (e-f) the milling motion remains a stable phase for all densities considered, with $P=0.015$, $M=0.998$ in Panel (e); $P=0.044$, $M=0.977$ in Panel (f); and $P=0.005$, $M=1.000$ in Panel (g). This shows the robustness of the local polarization dynamics driven by strong explicit alignment interactions.
Figure 8 presents the snapshots and MSD for simulations with a rough boundary. The corresponding temporal dynamics are displayed as videos in the Supplemental Material. Supplementary Movies 15-17 correspond to the snapshots in Fig. 8a-c and Supplementary Movies 18-20 corresponds to those in Fig. 8e-g. For small $R=0.03$, at low density $\phi=0.580$, we find again strong milling, with $P=0.017$, $M=0.990$, and an empty central region, as displayed in Panel (a). For higher densities, we observe increasing polarization in Panels (b) and (c), with $P=0.585$, $M=0.396$ for $\phi=0.907$, and $P=0.835,M=0.065$ for $\phi=1.306$. These are rotating polarized states formed by agents displaying localized rotation. For large $R=0.3$, as in the smooth boundary case, the milling state is much more robust and remains stable for all densities, with $P=0.023$, $M=0.996$ for $\phi=0.580$ in Panel (e); $P=0.107$, $M=0.933$ for $\phi=0.907$ in Panel (f); and $P=0.098$, $M=0.799$ for $\phi=1.306$ in Panel (g).
Discussion
We studied a minimal model of active polar disks with self-alignment in their dynamical equations and mutual alignment due to the the torques introduced by their off-centered rotation. We considered multiple model variations, including isotropic or anisotropic mobility, rough or smooth confinement, and different system sizes and densities. We observed a variety of emerging collective state with different structures and dynamics, combining polarized regions, milling, and vortex formation.
Despite the observed diversity of emerging states, we find that all display a combination of the slow rotational frequencies associated to milling and the high rotational frequencies of localized circular oscillations. The milling dynamics is mainly driven by mutual alignment forces that increase with $R$, resulting from a decentralized consensus process on the orientation that produces local polarization leading to collective rotation due to the circular confinement. The localized circular oscillations, on the other hand, are driven self-alignment and its tendency to excite low collective oscillatory modes, here corresponding to the combined horizontal and vertical oscillations that produce localized circular motion. In this framework, the various emergent states can be viewed as different superpositions of milling and local circular oscillations, which depend on $R$. These produce solid and fluid regimes in ordered and disordered configurations, depending on the repulsion versus self-propulsion forces and on the velocity differences between the various regions.
Our simulation results show that the proposed model for dense systems of active polar disks robustly produces a superposition of collective milling and localized oscillations for a variety of self-aligning dynamics, off-centered rotation distance $R$, isotropic or anisotropic disk-substrate damping, boundary conditions, etc. We thus note that the details of model implementation do not appear to affect the expected emerging collective states. Our proposed minimal model thus captures the generic combination of states that can be reached in dense active elastic systems, combining the effects of self-alignment and mutual alignment. It could therefore be used to describe a variety of real-world experimental biological and robotic actives systems.
Methods
Agent-based simulation timestep
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt$ (5)
$+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\cos\left(\theta_{i}\right)dt$
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt$ (6)
$+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt$
$+\mu_{\perp}\left[-f_{i}^{x}\sin\left(\theta_{i}\right)+f_{i}^{y}\cos\left(\theta_{i}\right)\right]\cos\left(\theta_{i}\right)dt\,.$
$\theta_{i}(t+dt)$ $=\theta_{i}\left(t\right)+\beta\left[-f_{i}^{x}\sin\left(\theta_{i}\right)+f_{i}^{y}\cos\left(\theta_{i}\right)\right]dt$ (7)
$+\sqrt{2D_{\theta}dt}~{}\eta_{i}(t)\,.$
We derive the equations for the isotropic and anisotropic limit cases considered in the paper. For isotropic disk-substrate damping, $\mu\equiv\mu_{\parallel}=\mu_{\perp}$, and equations (5) and (6) simplify to
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt+\mu f_{i}^{x}dt,$ (8)
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt+\mu f_{i}^{y}dt.$ (9)
For anisotropic damping, we study the extreme anisotropy case where $\mu_{\perp}=0$, with no possible displacement perpendicular to the orientation, as in the case of wheeled vehicles. In this case, equations (5) and (6) simplify to
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt$
$+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt,$ (10)
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt$
$+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt.$ (11)
Simulation arena
The simulation arena is defined by a smooth or rough circular confinement barrier, In the smooth circular confinement case, the outer ring boundary exerts a repulsive radial linear force inwards, perpendicular to its surface, as shown in Fig. 1b. In the rough circular confinement case, the boundary is composed of fixed disks with equivalent repulsive potential as the active disks in the simulation, as shown in Fig. 1c.
The radius of the circular arena is set to $R_{c}=10.5$ for $N=400$ soft polar disks, leading to a packing fraction $\phi=Nl_{0}^{2}/4R_{c}^{2}=0.907$, where $l_{0}=1$ is the disk diameter. We vary the radius of the circular confinement, $R_{c}$, for different total numbers of active polar disks, $N$, to maintain constant packing fraction, $\phi=Nl_{0}^{2}/4R_{c}^{2}$, as shown in Fig. 6. For $N=400,1600,6400$, we set $R_{c}=10.5,21,42$ with $l_{0}=1$, all corresponding to $\phi=0.907$.
For Figs. 7 and 8, we vary $l_{0}$ to adjust the density or packing fraction $\phi$. Specifically, we use $l_{0}=0.8,1.0,1.2$ to set the packing fraction to $\phi=0.580,0.907,1.306$ while keeping $R_{c}=10.5$ and $N=400$ fixed.
Simulation setup and parameters
All simulations are carried out starting from random initial positions and orientations. We consider homogeneous soft disks, all with the same diameter $l_{0}=1$ and self-propulsion speed $v_{0}=0.002$, unless otherwise noted. The force alignment strength coefficient is set to $\beta=1.2$ and the simulation time step is $dt=0.01$ in all simulations. We consider systems with $N=400$ agents in most of our analyses, while also performing simulations with $N=1600$ and $N=6400$ agents to study how the emerging states may change in larger systems.
We choose two values of the off-centered rotation distance $R$ to characterize, on one hand, a weak mutual alignment case with small $R=0.03$, and on the other, a strong mutual alignment case with large $R=0.3$. We also explore the $R$-dependent transition in vortex dynamics for multiple values of $R$, ranging from 00 to $0.3$ (see Supplementary Information).
We select two extreme cases of isotropic or anisotropic mobility: A fully isotropic case with $\mu\equiv\mu_{\parallel}=\mu_{\perp}$ and $\mu=0.02$, where the disk-substrate damping is equal in all directions, and a fully anisotropic damping case with $\mu_{\perp}=0$ and $\mu_{\parallel}=0.02$, where the agents can only move forward or backwards (as in the case of wheeled vehicles).
A detailed definition of all the simulation variables, along with their specific values in our simulations, are provided in Supplementary Tables 1 and 2.
Orientation autocorrelation
We compute the orientation temporal autocorrelation function $C(\tau)$ to evaluate how the particles’ $\hat{n}_{i}(t)$ orientations change over time and identify their rotational frequencies. We define this function for $N$ particles using the expression
Mean squared displacement
@@ Methods
The mean squared displacement (MSD) quantifies how far a particle has moved from its initial position over time, providing insights into its diffusive behavior within a system. The MSD for $N$ particles with position vectors $\vec{r}_{i}(t)$ is defined by [40]
@@
Vortex tracking
@@
Acknowledgments
We are grateful to Silke Henkes for helpful discussions. W.T. acknowledges the National Natural Science Foundation of China, Grant 62176022. Y.Z. and P.R. thank the support of Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-EXC 2002/1 ‘Science of Intelligence’, Project 390523135. The work of C.H. and A.S. was supported by the John Templeton Foundation, Grant 62213.
Author contributions
Competing interests
The authors declare no competing interests.
Data availability
The authors declare that the data supporting the findings of this study are available within the paper.
Code availability
The code to generate data using numerical simulations is available upon request.
Additional information
Supplementary information

### v3

@@ 3.5 State transitions as a function of density
To carry out our analyses, we compute the mean squared displacement (MSD) of the particle positions as a function of time for simulations with different densities and setups, to compare their corresponding dynamics based on the resulting curves. The mean squared displacement (MSD) quantifies how far a particle has moved from its initial position over time, providing insights into its diffusive behavior within a system. The MSD for $N$ particles with position vectors $\vec{r}_{i}(t)$ is defined by [40]
@@
We change the density by varying the disk diameter $l_{0}$ while keeping constant the number of agents $N=400$ and the size of the arena $R_{c}=10.5$. We study two types of boundary conditions, analyzing simulations in arenas with smooth or rough circular confinements, while focusing on anisotropic damping. Finally, we consider two scenarios for the off-centered rotation distance: a small $R=0.03$ case and a large $R=0.3$ case, corresponding to weak and strong explicit alignment, respectively. Note that $R$ is defined here as independent of disk diameter, so changing the value of $l_{0}$ will also change the relative off-centered rotation distance with respect to its radius. We thus find that, for $R=0.03$ and $l_{0}=0.8,1.0,1.2$, we have $R/(l_{0}/2)=0.075,0.06,0.05<<1$, all representing weak alignment. For $R=0.3$, we have $R_{r}=0.75,0.6,0.5$, which correspond instead to strong alignment.
Fig. 7 displays snapshots and the MSD for simulations in a smooth boundary arena, with different densities given by low ($\phi=0.580$) and high ($\phi=0.907$, $\phi=1.306$) packing fractions. The corresponding temporal dynamics are displayed as videos in the Supplemental Material. Supplementary Movies 9-11 correspond to the snapshots in Figs. 7a-c and Supplementary Movies 12-14 corresponds to the snapshots in Figs. 7e-g. For the small $R=0.03$ case, Panel (a) shows that, at low densities, we find clear milling motion around an empty center in a state characterized by order parameters $P=0.012$ and $M=0.997$. At higher densities, this milling state is suppressed as the central region is filled with disks and the order parameters become $P=0.387$ and $M=0.753$ in Panel (b) and $P=0.308$, $M=0.821$ in Panel (c), which corresponds more to an aligned state displaying localized rotation. On the other hand, for the large $R=0.3$ case in Panels (e-f) the milling motion remains a stable phase for all densities considered, with $P=0.015$, $M=0.998$ in Panel (e); $P=0.044$, $M=0.977$ in Panel (f); and $P=0.005$, $M=1.000$ in Panel (g). This shows the robustness of the local polarization dynamics driven by strong explicit alignment interactions.
Fig. 8 presents the snapshots and MSD for simulations with a rough boundary. The corresponding temporal dynamics are displayed as videos in the Supplemental Material. Supplementary Movies 15-17 correspond to the snapshots in Figs. 8a-c and Supplementary Movies 18-20 corresponds to those in Figs. 8e-g. For small $R=0.03$, at low density $\phi=0.580$, we find again strong milling, with $P=0.017$, $M=0.990$, and an empty central region, as displayed in Panel (a). For higher densities, we observe increasing polarization in Panels (b) and (c), with $P=0.585$, $M=0.396$ for $\phi=0.907$, and $P=0.835,M=0.065$ for $\phi=1.306$. These are rotating polarized states formed by agents displaying localized rotation. For large $R=0.3$, as in the smooth boundary case, the milling state is much more robust and remains stable for all densities, with $P=0.023$, $M=0.996$ for $\phi=0.580$ in Panel (e); $P=0.107$, $M=0.933$ for $\phi=0.907$ in Panel (f); and $P=0.098$, $M=0.799$ for $\phi=1.306$ in Panel (g).
4 Conclusion
We studied a minimal model of active polar disks with self-alignment in their dynamical equations and mutual alignment resulting from torques introduced by their off-centered rotation. We considered multiple model variations, including isotropic or anisotropic mobility, rough or smooth confinement, and different system sizes and densities. We identified a variety of emerging collective state with different structures and dynamics, combining polarized regions, milling, and vortex formation.
Despite the observed diversity of emerging states, they all display a combination of slow rotational frequencies associated to milling and high rotational frequencies linked to localized circular oscillations. The milling dynamics is mainly driven by mutual alignment forces that increase with $R$, resulting from a decentralized consensus process on the orientation that produces local polarization leading to collective rotation due to the circular confinement. The localized circular oscillations, on the other hand, are driven by self-alignment and its tendency to excite low collective oscillatory modes, here corresponding to the combined horizontal and vertical oscillations that produce localized circular motion. In this framework, the different emergent states reflect the combined effect of self- and mutual alignment as a function of $R$, which can produce solid or fluid regimes in ordered or disordered configurations, depending on the specific setup and parameters. Their analysis thus opens the door to a better understanding of these two self-organizing mechanisms in active matter.
Our simulations demonstrate that the mechanical model of a dense system of active polar disks robustly produces a superposition of collective milling and localized oscillations across various scenarios, including self-aligning dynamics, off-centered rotation distances, isotropic or anisotropic disk-substrate damping, and boundary conditions. The implementation details of the model appear not to affect the emergence of these collective states. Thus, this minimal model captures the generic combination of states that can be realized in dense active elastic matter, integrating the effects of self-alignment and mutual alignment. It could therefore be used to describe a variety of real-world biological and robotic active systems.
Acknowledgements
We are grateful to Silke Henkes for helpful discussions.
Author contributions
Funding information
W.T. acknowledges the National Natural Science Foundation of China, Grant 62176022. Y.Z. and P.R. thank the support of Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-EXC 2002/1 ‘Science of Intelligence’(SCIoI), Project 390523135. The work of C.H. and A.S. was supported by the John Templeton Foundation, Grant 62213.
Appendix A Agent-based Simulation Timestep
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\cos\left(\theta_{i}\right)dt$ (A.1)
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt$ (A.2)
$+\mu_{\perp}\left[-f_{i}^{x}\sin\left(\theta_{i}\right)+f_{i}^{y}\cos\left(\theta_{i}\right)\right]\cos\left(\theta_{i}\right)dt,$
$\theta_{i}(t+dt)=\theta_{i}\left(t\right)+\beta\left[-f_{i}^{x}\sin\left(\theta_{i}\right)+f_{i}^{y}\cos\left(\theta_{i}\right)\right]dt+\sqrt{2D_{\theta}dt}\eta_{i}(t).$ (A.3)
We derive the equations for the isotropic and anisotropic limit cases considered in the paper. For isotropic disk-substrate damping, $\mu\equiv\mu_{\parallel}=\mu_{\perp}$, and equations (A.1) and (A.2) simplify to
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt+\mu f_{i}^{x}dt,$ (A.4)
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt+\mu f_{i}^{y}dt.$ (A.5)
For anisotropic damping, we study the extreme anisotropy case where $\mu_{\perp}=0$, with no possible displacement perpendicular to the orientation, as in the case of wheeled vehicles. In this case, equations (A.1) and (A.2) simplify to
$x_{i}(t+dt)$ $=x_{i}\left(t\right)+v_{0}\cos\left(\theta_{i}\right)dt+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt,$ (A.6)
$y_{i}(t+dt)$ $=y_{i}\left(t\right)+v_{0}\sin\left(\theta_{i}\right)dt+\mu_{\parallel}\left[f_{i}^{x}\cos\left(\theta_{i}\right)+f_{i}^{y}\sin\left(\theta_{i}\right)\right]\sin\left(\theta_{i}\right)dt.$ (A.7)
A summary of all the simulation variables and their corresponding values is provided in Table 1. All our results are provided in natural units defined by the equations of motion and the parameter values provided above. For example, since we defined $l_{0}=1$, the unit of length in all our results is given by the disk diameter.
Table 1: Parameters and Default Values
Parameter Default value Definition
$N$ 400, 1600, 6400 Total number of particles
$l_{0}$ 1 Disk diameter
$R_{c}$ 10.5 Radius of the arena
$\phi=Nl_{0}^{2}/4R_{c}^{2}$ $0.907$ Packing fraction
$v_{0}$ 0.002 Active speed
[…]
$k$ 5 Elastic coefficient
$\beta$ 1.2 Force alignment strength
[…]
$\mu_{\parallel}$ 0.02 Parallel mobility coefficient
$\mu_{\perp}$ 0.00 or 0.02 Perpendicular mobility coefficient
$dt$ 0.01 Simulation time step
$T$ $2\times 10^{7}$ Total number of time steps
$T\times dt$ $2\times 10^{5}$ Total simulation time
$n_{\mathrm{rep}}$ 5 Number of simulation repetitions
Appendix B Phase Diagrams
[…]
Figure 10: Polarization and milling as a function of control parameter R. Polarization $P$ and milling order parameter $M$ as a function of $R$ in systems with isotropic or anisotropic agent-substrate damping and with a rough or smooth boundary. Here, $l_{0}=1.0$, $v_{0}=0.002$, and each point is averaged over $20$ realizations.
Figure 9 presents exploratory phase diagrams in a two-dimensional parameter space, as a function of the off-centered rotation distance $R$ and noise $D_{\theta}$. We display the polarization $P$ and milling order parameter $M$ as heat maps for the four combinations of isotropic or anisotropic disk-substrate interactions, and of smooth or rough boundaries, that we consider in this work. In addition to these diagrams, Fig. 10 plots the values of $P$ and $M$ as a function of $R$ for low and intermediate noise levels.
These figures show that, for a range of low noise levels, we typically find either regions with low $P$ and high $M$, corresponding to the milling regime, or with high $P$ and low $M$, corresponding to localized rotation. This suggests that these states can emerge in real-world systems with sufficiently low levels of noise, which are the focus of this work. For high noise values, we instead mainly find states with low $P$ and low $M$ that display either quenched or dynamic disorder [55].
Appendix C Vortex Tracking
@@
Appendix D Results in Large Systems
Fig. 12 presents the orientation autocorrelation functions for the same parameter combinations presented in Fig. 3 of the main paper, but for a larger system with $N=1600$. Fig. 13 displays the Fourier transforms of these correlation functions, corresponding to the same cases shown in Fig. 4 of the main article, but for a larger $N=1600$.
Both figures show that the low- and high-frequency oscillations observed in each case are equivalent to those described in the paper for smaller systems.
Fig. 15 display the frequency and amplitude of the main low- and high-frequency peaks as a function of $R$ for different combinations of isotropic (I) or anisotropic (A) agent-substrate interactions and a smooth (S) or rough (R) confining boundary, as indicated by the column titles. Fig. 15 displays large system case with $N=1600$, keeping the same mean density or packing fraction with Fig. 14.
The figures show that, the high-frequency circular oscillations can always be found at small values of $R$, while the low-frequency milling appears at all $R$ values. This is true for all simulation conditions and for both considered system sizes, $N=400$ and $N=1600$.
Figure 12: Orientation temporal autocorrelation functions. The displayed plots are equivalent to those presented in Fig. 3 of the main paper, but here for a larger system with $N=1600$ and the same packing fraction. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0$.
Figure 13: Fourier transforms of orientation autocorrelation functions. The displayed plots are equivalent to those presented in Fig. 4 of the main paper, for a larger system with $N=1600$ and the same packing fraction. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0$.
Figure 14: Frequencies and amplitudes of orientation autocorrelation peaks for $N=400$. The low-frequency (LF) and high-frequency (HF) oscillation peaks are identified in the Fourier transform of the orientation autocorrelation function, and their frequency and amplitude is displayed as a function of the off-centered rotation distance $R$. The frequency and amplitude plots are displayed for four different cases of isotropic (I) or anisotropic (A) agent-substrate mobility and smooth (S) or rough (R) boundary. We present the I/S case in panels a,e; the I/R case in b,f; the A/S case in c,g; and the A/R case in d,h. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0$.
Figure 15: Frequencies and amplitudes of orientation autocorrelation peaks for $N=1600$. The low-frequency (LF) and high-frequency (HF) oscillation peaks are identified in the Fourier transform of the orientation autocorrelation function, and their frequency and amplitude is displayed as a function of the off-centered rotation distance $R$. The frequency and amplitude plots are displayed for four different cases of isotropic (I) or anisotropic (A) agent-substrate mobility and smooth (S) or rough (R) boundary. We present the I/S case in panels a,e; the I/R case in b,f; the A/S case in c,g; and the A/R case in d,h. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0$.
Appendix E Vortex Dynamics
In the next four figures, we present scatter plots of the locations where vortices appear during a simulation run, for cases with isotropic or anisotropic mobility and smooth or rough circular confinement. In each figure the identified vortex positions are displayed as red dots for different off-centered distances $R=0.0$, $0.03$, $0.3$, and different system sizes $N=400$, $1600$, $6400$, while keeping the packing fraction $\phi$ constant. The number indicated in each panel represents the mean number of vortices identified per snapshot. The cases where this number is approximately equal to $1$ typically display either a single milling vortex at the center of the arena or a single orbiting vortex separating the outer milling and inner oscillating regions.
More specifically, Fig. 16 shows cases with isotropic mobility and a smooth confining boundary, Fig. 17 shows cases with isotropic mobility and a rough confinement, Fig. 18 shows cases with anisotropic mobility and smooth confinement, and Fig. 19 shows cases with anisotropic mobility and rough confinement.
We observe, in general, that larger systems tend to have more vortices, but with less structured positions and trajectories. In the smooth boundary case, vortices form mainly in the bulk. In the localized rotation regime, they form a ring of vortices that matches the inner and outer global vorticity in small systems while moving throughout the arena in large systems. In the rough boundary case, vortices mainly nucleate near the edge of the system and then move into the bulk. Finally, we note that for larger system sizes the vortex nucleation and spatial dynamics become increasingly chaotic.
Figure 16: Vortex positions for isotropic mobility and a smooth boundary. The red points indicate the vortex positions identified during a run for different off-centered rotation distances and system sizes, keeping the density constant. The number in each panel displays the mean number of vortices per snapshot. Here, agent-substrate interactions are isotropic and the confining boundary is smooth. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.0$.
Figure 17: Vortex positions for isotropic mobility and a rough boundary. The red points indicate the vortex positions identified during a run for different off-centered rotation distances and system sizes, keeping the density constant. The number in each panel displays the mean number of vortices per snapshot. Here, agent-substrate interactions are isotropic and the confining boundary is rough. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.0$.
Figure 18: Vortex positions for anisotropic mobility and a smooth boundary. The red points indicate the vortex positions identified during a run for different off-centered rotation distances and system sizes, keeping the density constant. The number in each panel displays the mean number of vortices per snapshot. Here, agent-substrate interactions are anisotropic and the confining boundary is smooth. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.0$.
Figure 19: Vortex positions for anisotropic mobility and a rough boundary. The red points indicate the vortex positions identified during a run for different off-centered rotation distances and system sizes, keeping the density constant. The number in each panel displays the mean number of vortices per snapshot. Here, agent-substrate interactions are anisotropic and the confining boundary is rough. Other simulation parameters are set to: $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.0$.